OpenStudy (anonymous):

300 million test kits are sold this year, if there is 115% increase from last year, how many test kits were sold last year?

OpenStudy (jamesj):

Suppose X were sold last year and Y this year. Then the number \[ \frac{Y - X}{X} = \frac{Y}{X} - 1 \] is the percentage increase. Now if the number has increased by 115% that means that number is 1.15. You know the number X. Solve for Y.

OpenStudy (jamesj):

\[ \frac{Y}{X} - 1 = 1.15 \] or \[ \frac{Y}{X} = 2.25 \]

OpenStudy (jamesj):

*correction: you know the number Y, the number sold THIS year. You want to solve for X, the number sold last year.

OpenStudy (jamesj):

Make sense?

OpenStudy (campbell_st):

just work with 115% = 300000 the 1% = 300000/115 then multiply by 100 to get the correct answer 260 870.

OpenStudy (jamesj):

That's not the correct answer.

OpenStudy (campbell_st):

quick question... how does X/Y - 1 = 1.15 then become X/Y = 2.25... I think that gives a wrong answer...

OpenStudy (campbell_st):

and geee...add 3 zeros to my answer... to get it correct